# Gravity from thermodynamics in vacuum: Lorentz invariance and the role of Bel-Robinson tensor

### v1

@@ III Higher order corrections in vacuum: the Bel-Robinson tensor
$\sigma_{\mu\nu}\left(x\right)=$ $\lambda\left[C_{\lambda\rho\sigma\tau}k^{\lambda}k^{\sigma}h_{\mu}^{\rho}h_{\nu}^{\tau}-+\frac{1}{2}\left(h_{\mu\lambda}h_{\nu\rho}-\frac{1}{2}h_{\mu\nu}h_{\lambda\rho}\right)R^{\lambda\rho}\right]\left(-1+\frac{\lambda\theta_{\text{flat}}}{1+\lambda\theta_{\text{flat}}}\right)$
@@
$\int m^{\mu}m^{\nu}m^{\lambda}m^{\rho}\text{d}\Omega_{2}=$ $\frac{4\pi}{15}\left(\gamma^{\mu\nu}\gamma^{\lambda\rho}+\gamma^{\mu\lambda}\gamma^{\nu\rho}+\gamma^{\mu\rho}\gamma^{\nu\lambda}\right),$ (30)
where we introduced the spatial metric $\gamma^{\mu\nu}=g^{\mu\nu}+n^{\mu}n^{\nu}$. The integration in the null parameter $\lambda$ is straightforward. In the end, we arrive at the following expression for $\Delta\mathcal{A}_{\text{rev}}$
@@
where we discarded the $O\left(l^{6}\right)$ terms [^3] Justifying the removal of these terms is in fact rather subtle. To treat the spacetime as a differentiable manifold, we must keep the size parameter of the causal diamond much larger than the Planck scale. Then, $O\left(l^{6}\right)$ contribution can actually be larger than the $O\left(l^{4}l_{\text{P}}^{2}\right)$ ones. However, the size parameter $l$ is only restricted to be much larger than the Planck length and at the same time much smaller than the local curvature length scale. Within this range, the value of $l$ is completely arbitrary. Therefore, at any regular spacetime point, we can construct a sequence of causal diamonds with different size parameters. For any such diamond, we can derive a version of equation (33), differing only in the value of $l$. Consequently, equation (33) is only satisfied if the terms proportional to different powers of $l$ vanish separately. This observation allows us to focus on the leading order $O\left(l^{4}\right)$ part of this equation and disregard all the higher order contributions.. Since the Ricci tensor is $O\left(l_{\text{P}}^{2}\right)$, the second term only contributes at $O\left(l_{\text{P}}^{4}\right)$ and we can also discard it, leaving the following equation
@@ IV Interpretation of the Bel-Robinson tensor contribution
We start by briefly reviewing the properties of the Bel-Robinson tensor $T_{\mu\nu\lambda\rho}$, which appears as the source term in equation (35). The Bel-Robinson tensor in $4$ spacetime dimensions (the only case we consider) is given by equation (32). It is symmetric in all of its indices and traceless (these properties are specific to $4$ dimensions). In any spacetime dimension, it obeys an analogue of the dominant energy condition, i.e., for any four timelike, future-pointing vector $v_{i}^{\mu}$, $i\in\left[1,4\right]$, it holds $T_{\mu\nu\lambda\rho}v_{1}^{\mu}v_{2}^{\nu}v_{3}^{\lambda}v_{4}^{\rho}\geq 0$ and the vector $-T_{\mu\nu\lambda\rho}v_{1}^{\nu}v_{1}^{\lambda}v_{1}^{\rho}$ is future-pointing. The Bel-Robinson tensor is non-degenerate, in the sense that it vanishes if and only if the Weyl tensor is equal to zero. Moreover, its divergence is identically zero in the Einstein spacetimes (i.e., the spacetimes with constant Ricci curvature). The Bel-Robinson tensor is in fact the unique tensor quadratic in the Weyl tensor with these properties [50, 70, 71].
@@
In our case, the Bel-Robinson super-energy density multiplied by a Planck scale constant with the dimensions of area $-2\mathcal{C}l_{\text{P}}^{2}/\left(25\pi\right)$ appears as a source for the equations governing the gravitational dynamics. Therefore, it is tempting to interpret our result as the modified Einstein equations in which the Ricci curvature is sourced by the quasilocal gravitational energy density integrated over some Planck-scale unit of area. However, as we show in the following, embracing this interpretation requires that we surrender the local Lorentz invariance of physics at the Planck scale.
@@ IV.1 Locally Lorentz invariant interpretation
In A we show that this equation implies that the fully completely symmetrised part of the tensor vanishes
@@
$08\left(R_{\mu\nu}-\frac{1}{4}Rg_{\mu\nu}\right)+6\frac{2\mathcal{C}l_{\text{P}}^{2}}{25\pi}T_{\mu\nu\lambda}^{\;\>\;\>\;\>\lambda}=$ 0 . (41)
@@
In other words, thermodynamic reasoning leads to no local vacuum corrections to gravitational dynamics at the order $O\left(l_{\text{P}}^{2}\right)$ if the local Lorentz invariance is preserved. This is consistent with the non-existence of any local, purely metric, diffeomorphism-invariant (or WTDiff-invariant) theory in $4$ spacetime dimensions whose action is at most quadratic in the Riemann tensor whose equations of motion are not solved by the vacuum solutions of general relativity [74] (or Weyl transverse gravity [75]).
@@ IV.3 Conceptual reflections on the possible Lorentz invariance violations
  • Both the local causal horizons and the Unruh effect implicitly rely on the local Lorentz invariance. Nevertheless, both concepts have been studied in the literature in some local Lorentz invariance violating settings and found to be robust even in that case [80, 81, 82]. One would just need to carefully check their validity in the specific scenario we consider.
@@
  • Further analysis of the local Lorentz invariance interpretation of the local equilibrium conditions (especially going to the sixth order in the derivatives of the metric) might lead to a result more similar to Hořava-Lifshitz gravity, possibly even corresponding to some specific version of that theory. Given the status of Hořava-Lifshitz gravity as a possible UV completion of general relativity, obtaining it from thermodynamics of local causal horizons would be quite interesting.
@@ V Discussion
In this work, we examine whether local equilibrium conditions encode corrections to the traceless Einstein equations in vacuum. In previous works, we have shown that perturbative corrections of this kind do occur in the presence of matter fields [60, 61, 62]. However, these works focused on extracting specific types of correction terms and neglected the possible contributions quadratic in Weyl tensor on which we focus here. Our main message is that, under the assumptions of locality and local Lorentz invariance, local equilibrium conditions imply no modifications to the equations governing gravitational dynamics that are perturbatively relevant in the vacuum at the order $O\left(l_{\text{P}}^{2}\right)$. In other words, the vacuum solutions to the Einstein equations acquire no corrections in this regime (although there might exist novel non-perturbative vacuum solutions).
@@
  • We consider the logarithmic term in entropy of local causal horizons. The leading order term proportional to area is $O\left(1/l_{\text{P}}^{2}\right)$, whereas the logarithmic correction is $O\left(l_{\text{P}}^{0}\right)$. Thence, including this term bring a new length scale in the form of the Planck length, which makes the presence of local corrections to the Einstein equations possible (since such corrections have necessarily different dimension than the Ricci tensor).
@@
We also proposed an alternative, local Lorentz invariance violating, interpretation, which does imply modified gravitational dynamics. The result has the form of the Bel-Robinson super-energy density multiplied by a Planck scale area acting as a source for the time-time component of the traceless Ricci tensor. This outcome agrees remarkably well with the interpretation of the Bel-Robinson super-energy density as the quasilocal gravitational energy density per unit area.
@@ Appendix A Removing contractions with an arbitrary timelike vector
In this appendix, we show that if a rank $k$ tensor $Q_{\alpha_{1}\dots\alpha_{k}}$ obeys $Q_{\alpha_{1}\dots\alpha_{k}}n^{\alpha_{1}}\dots n^{\alpha_{k}}=0$ for an arbitrary unit, timelike, future-pointing vector $n^{\mu}$, its fully symmetric part must vanish. For this purpose, we introduce an orthonormal basis $e^{0}$, $e^{1}$,…,$e^{D-1}$. In this basis, any unit, timelike, future-pointing vector $n^{\mu}$ can be decomposed as
@@ Appendix B Non-local results for general spacetime dimensions
In the main text, we strictly assumed the equations governing gravitational dynamics to be local. Then, these equations cannot depend on an arbitrary diamond size parameter $l$. At this point, we cannot physically motivate any scenario where $l$ does affect the gravitational dynamics. Nevertheless, the $l$-dependent results are easy to obtain and rather interesting in their own right. Hence, we briefly discuss them in this appendix. Let us then assume that $l$ either takes a preferred value (corresponding to some non-locality scale), or that our choice of $l$ somehow affects the gravitational dynamics. We still require the condition $l\gg l_{\text{P}}$ to view the spacetime as a smooth manifold. Consequently, the $l$-dependent corrections to the equations become much larger than the corrections coming from the logarithmic contribution to entropy, which are suppressed by $l_{\text{P}}^{2}$, and we can work with a simple vacuum equilibrium condition $\Delta\mathcal{A}_{\text{rev}}=0$. Since we do not rely on the logarithmic term which is the leading order correction to entropy only in $4$ spacetime dimensions, we can carry out the derivation in an arbitrary dimension $D$. In this case, we find for $\Delta\mathcal{A}_{\text{rev}}$

### v2

@@ III Higher order corrections in vacuum: the Bel-Robinson tensor
$\sigma_{\mu\nu}\left(x\right)=$ $\lambda\left[C_{\lambda\rho\sigma\tau}k^{\lambda}k^{\sigma}h_{\mu}^{\rho}h_{\nu}^{\tau}-\frac{1}{2}\left(h_{\mu\lambda}h_{\nu\rho}-\frac{1}{2}h_{\mu\nu}h_{\lambda\rho}\right)R^{\lambda\rho}\right]\left(-1+\frac{\lambda\theta_{\text{flat}}}{1+\lambda\theta_{\text{flat}}}\right)$
@@
$\int m^{\mu}m^{\nu}m^{\lambda}m^{\rho}\text{d}\Omega_{2}=$ $\frac{4\pi}{15}\left(\gamma^{\mu\nu}\gamma^{\lambda\rho}+\gamma^{\mu\lambda}\gamma^{\nu\rho}+\gamma^{\mu\rho}\gamma^{\nu\lambda}\right).$ (32)
The integration in the null parameter $\lambda$ is straightforward. In the end, we arrive at the following expression for $\Delta\mathcal{A}_{\text{rev}}$
@@
where we discarded the $O\left(l^{6}\right)$ terms [^5] Justifying the removal of these terms is in fact rather subtle. To treat the spacetime as a differentiable manifold, we must keep the size parameter of the causal diamond much larger than the Planck scale. Then, $O\left(l^{6}\right)$ contribution can actually be larger than the $O\left(l^{4}l_{\text{P}}^{2}\right)$ ones. However, the size parameter $l$ is only restricted to be much larger than the Planck length and at the same time much smaller than the local curvature length scale. Within this range, the value of $l$ is completely arbitrary. Therefore, at any regular spacetime point, we can construct a sequence of causal diamonds with different size parameters. For any such diamond, we can derive a version of equation (35), differing only in the value of $l$. Consequently, equation (35) is only satisfied if the terms proportional to different powers of $l$ vanish separately. This observation allows us to focus on the leading order $O\left(l^{4}\right)$ part of this equation and disregard all the higher order contributions. For the details of this procedure see a previous work of the authors [82].. Since the Ricci tensor is $O\left(l_{\text{P}}^{2}\right)$, the second term only contributes at $O\left(l_{\text{P}}^{4}\right)$ and we can also discard it, leaving the following equation
@@ IV Interpretation of the Bel-Robinson tensor contribution
We start by briefly reviewing the properties of the Bel-Robinson tensor $T_{\mu\nu\lambda\rho}$, which appears as the source term in equation (37). The Bel-Robinson tensor in $4$ spacetime dimensions (the only case we consider) is given by equation (34). It is symmetric in all of its indices and traceless (these properties are specific to $4$ dimensions). In any spacetime dimension, it obeys an analogue of the dominant energy condition, i.e., for any four timelike, future-pointing vector $v_{i}^{\mu}$, $i\in\left[1,4\right]$, it holds $T_{\mu\nu\lambda\rho}v_{1}^{\mu}v_{2}^{\nu}v_{3}^{\lambda}v_{4}^{\rho}\geq 0$ and the vector $T_{\mu\nu\lambda\rho}v_{1}^{\nu}v_{1}^{\lambda}v_{1}^{\rho}$ is future-pointing. The Bel-Robinson tensor is non-degenerate, in the sense that it vanishes if and only if the Weyl tensor is equal to zero. Moreover, its divergence is identically zero in the Einstein spacetimes (i.e., the spacetimes with constant Ricci curvature). The Bel-Robinson tensor is in fact the unique tensor quadratic in the Weyl tensor with these properties [50, 94, 95].
@@
In our case, the Bel-Robinson super-energy density multiplied by a Planck scale constant with the dimensions of area $2\mathcal{C}l_{\text{P}}^{2}/\left(25\pi\right)$ appears as a source for the equations governing the gravitational dynamics. Therefore, it is tempting to interpret our result as the modified Einstein equations in which the Ricci curvature is sourced by the quasilocal gravitational energy density integrated over some Planck-scale unit of area. However, as we show in the following, embracing this interpretation requires that we surrender the local Lorentz invariance of physics at the Planck scale.
@@ IV.1 Locally Lorentz invariant interpretation
In A we show that this equation implies vanishing of the fully symmetrised part of the tensor
@@
$8\left(R_{\mu\nu}-\frac{1}{4}Rg_{\mu\nu}\right)+6\frac{2\mathcal{C}l_{\text{P}}^{2}}{25\pi}T_{\mu\nu\lambda}^{\;\>\;\>\;\>\lambda}=0.$ (43)
@@
In other words, thermodynamic reasoning leads to no local vacuum corrections to gravitational dynamics at the order $O\left(l_{\text{P}}^{2}\right)$ if the local Lorentz invariance is preserved. This is consistent with the non-existence of any local, purely metric, diffeomorphism-invariant (or WTDiff-invariant) theory in $4$ spacetime dimensions whose action is at most quadratic in the Riemann tensor and whose equations of motion are not solved by the vacuum solutions of general relativity [98] (or Weyl transverse gravity [71]).
@@ IV.3 Conceptual reflections on the possible Lorentz invariance violations
  • Both the local causal horizons and the Unruh effect implicitly rely on the local Lorentz invariance. However, both concepts have been studied in the literature in some local Lorentz invariance violating settings and found to be robust even in that case [103, 104, 105]. One would just need to carefully check their validity in the specific scenario we consider.
@@
  • Further analysis of the local Lorentz invariance violating interpretation of the local equilibrium conditions (especially going to the sixth order in the derivatives of the metric) might lead to a result more similar to Hořava-Lifshitz gravity, possibly even corresponding to some specific version of that theory. Given the status of Hořava-Lifshitz gravity as a possible UV completion of general relativity, obtaining it from thermodynamics of local causal horizons would be quite interesting.
@@ V Discussion
In this work, we examine whether local equilibrium conditions encode corrections to the traceless Einstein equations in vacuum. Previously, we have shown that perturbative corrections of this kind do occur in the presence of matter fields [82, 83, 84]. However, these works focused on extracting specific types of correction terms and neglected the possible contributions quadratic in Weyl tensor on which we focus here. Our main message is that, under the assumptions of locality and local Lorentz invariance, local equilibrium conditions imply no modifications to the equations governing gravitational dynamics that are perturbatively relevant in the vacuum at the order $O\left(l_{\text{P}}^{2}\right)$. In other words, the vacuum solutions to the Einstein equations acquire no corrections in this regime (although there might exist novel non-perturbative vacuum solutions).
@@
  • We consider the logarithmic term in entropy of local causal horizons. The leading order term proportional to area is $O\left(1/l_{\text{P}}^{2}\right)$, whereas the logarithmic correction is $O\left(l_{\text{P}}^{0}\right)$. Thence, including this term bring a new length scale in the form of the Planck length, which makes the presence of local corrections to the Einstein equations possible (since such corrections have necessarily different length dimension than the Ricci tensor).
@@
We also proposed an alternative, local Lorentz invariance violating, interpretation, which does imply modified gravitational dynamics. The result has the form of the Bel-Robinson super-energy density multiplied by a Planck scale area acting as a source for the time-time component of the traceless Ricci tensor. This outcome agrees remarkably well with the interpretation of the Bel-Robinson super-energy density as the quasilocal gravitational energy density per unit area. We stress that this alternative approach faces a number of challenges, most notably in recovering its the thermodynamic assumptions without the local Lorentz invariance. Moreover, we have been able to only outline it, as we presently lack the remaining equations of motion. Despite its shortcomings, this approach also has a number of attractive features. Due to the natural appearance of the Bel-Robinson tensor, it implies perturbative modifications to vacuum gravitational dynamics. Furthermore, it offers the possibility of making contact with the power-counting renormalisable Hořava-Lifshitz gravity. In conclusion, we believe that the Lorentz invariance violating interpretation of our results deserves further attention.
@@ Appendix A Removing contractions with an arbitrary timelike vector
In this appendix, we show that if a rank $k$ tensor $Q_{\alpha_{1}\dots\alpha_{k}}$ obeys $Q_{\alpha_{1}\dots\alpha_{k}}n^{\alpha_{1}}\dots n^{\alpha_{k}}=0$ for an arbitrary unit, timelike, future-pointing vector $n^{\mu}$, its fully symmetric part must vanish. For this purpose, we introduce an orthonormal basis $e^{0}$, $e^{1}$$e^{D-1}$. In this basis, any unit, timelike, future-pointing vector $n^{\mu}$ can be decomposed as
@@ Appendix B Non-local results for general spacetime dimensions
In the main text, we strictly assumed the equations governing gravitational dynamics to be local. Then, these equations cannot depend on an arbitrary diamond size parameter $l$. At this point, we cannot physically motivate any scenario where $l$ does affect the gravitational dynamics. Nevertheless, the $l$-dependent results are easy to obtain and rather interesting in their own right. Hence, we briefly discuss them in this appendix. Let us then assume that $l$ either takes a preferred value (corresponding to some non-locality scale), or that our choice of $l$ somehow affects the gravitational dynamics. We still require the condition $l\gg l_{\text{P}}$ to view the spacetime as a smooth manifold. Consequently, the $l$-dependent corrections to the equations become much larger than the corrections coming from the logarithmic contribution to entropy, which are suppressed by $l_{\text{P}}^{2}$, and we can work with a simple vacuum equilibrium condition $\Delta\mathcal{A}_{\text{rev}}=0$. Since we do not rely on the logarithmic term, which is the leading order correction to entropy only in $4$ spacetime dimensions, we can carry out the derivation in an arbitrary dimension $D$. In this case, we find for $\Delta\mathcal{A}_{\text{rev}}$